OpenStudy (anonymous):

the quantity of x plus 4 all over 2 = 7.

OpenStudy (anonymous):

i really need help!!

OpenStudy (marissalovescats):

Hi and welcome to OpenStudy!! Is this your problem and you want to solve for x? 

OpenStudy (anonymous):

First multiply 2 on both sides.

OpenStudy (jmark):

Given (x+4)/2 = 7 x+4 = 7x2 x+4 = 14 x = 14-4 x = 10 ICSE Class 8 English Syllabus - http://edworld.hpage.co.in/icse-class-8-english-syllabus_12856532.html